OpenStudy (jiteshmeghwal9):

If \[x=\sqrt{3+\sqrt{3+\sqrt{3+....}}}\], then,

OpenStudy (lgbasallote):

\[x^2 = 3 + x\]

OpenStudy (lgbasallote):

now solve for x

OpenStudy (jiteshmeghwal9):

\[Ans. 2<x <3\]

OpenStudy (jiteshmeghwal9):

I tried this but i don't get it:(

OpenStudy (lgbasallote):

do you get \[x^2 = 3+x\]

OpenStudy (jiteshmeghwal9):

Ya!

OpenStudy (lgbasallote):

\[x^2 - x - 3 = 0\] use Q.F> \[x = \frac{-(-1) \pm \sqrt{(-1)^2 - 4(1)(-3)}}{2(1)}\] \[x = \frac{1 \pm \sqrt{1 + 12}}{2}\] \[x = \frac{1 \pm \sqrt{13}}{2}\]

OpenStudy (lgbasallote):

then get the positive answer

OpenStudy (lgbasallote):

you will get something around \[x = 2.3\] got it?

OpenStudy (lgbasallote):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

No @lgbasallote

OpenStudy (lgbasallote):

which part

OpenStudy (jiteshmeghwal9):

Plz give me full solution:)

OpenStudy (lgbasallote):

maybe from the start... \[x = \sqrt{3+\sqrt{3+ \sqrt{ 3+ \cdots}}}\] \[x^2 = 3 + \sqrt{3+\sqrt{3 +\cdots}}\] \[x^2 = 3+x\] \[x^2 - x-3 = 0\] you need to solve for x so use quadratic formula

OpenStudy (lgbasallote):

better?

OpenStudy (jiteshmeghwal9):

U mean \[\alpha={1+\sqrt{1-12} \over 2}, \beta={1-\sqrt{-11} \over 2} \]

OpenStudy (lgbasallote):

-4(1)(-3) = +12

OpenStudy (lgbasallote):

negative times negative = postive remember?

OpenStudy (jiteshmeghwal9):

yes, but how we can gt \[2<x <3\]

OpenStudy (lgbasallote):

\[\frac{1+\sqrt{13}}{2} = 2.3\]

OpenStudy (lgbasallote):

try it in your calculator

OpenStudy (lgbasallote):

\(\sqrt{13} = 3.6\)

OpenStudy (anonymous):

lgba is right, this is concerned with quadratic formula.

OpenStudy (lgbasallote):

1+ 3.6 = 4.6 4.6/2 = 2.3

OpenStudy (jiteshmeghwal9):

I gt 3.16 (approximately).

OpenStudy (jiteshmeghwal9):

sorry, 3.60

OpenStudy (jiteshmeghwal9):

@lgbasallote

OpenStudy (lgbasallote):

that's sqrt 13 right?

OpenStudy (jiteshmeghwal9):

ya

OpenStudy (lgbasallote):

so you have \[\frac{1+\sqrt{13}}{2} \implies \frac{1+3.6}{2} \implies \frac{4.6}{2} = 2.3\]

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (lgbasallote):

\[x =\sqrt{3+\sqrt{3+\sqrt{3+\cdots}}}\] \[x^2 = 3 + \sqrt{3+\sqrt{3+\cdots}}\] \[x^2 = 3 + x\] \[x^2 - x - 3 = 0\] \[x = \frac{-(-1) \pm \sqrt{(-1) - 4(1)(-3)}}{2(1)}\] \[x = \frac{1 \pm \sqrt{1 + 12}}{2}\] \[x = \frac{1\pm \sqrt{13}}{2}\] \[x = 2.3\]

OpenStudy (jiteshmeghwal9):

K ! thanx a lot @lgbasallote